# Float Self-Tagging

### v2

@@ 6.2. Comparison to NaN-Boxing and NuN-Boxing
Figure 11. Execution time comparison between Bigloo with 1-tag self-tagging, Bigloo with NuN-boxing, and Bigloo with NaN-boxing. The y-axis shows execution time relative to the original Bigloo version, which uses tagged pointers. 1.0 indicates no change, lower means faster execution than the original version.
6.3. Application to JavaScript
This section experiments with the application of float self-tagging in JavaScript, which is a prime candidate for self-tagging because it has no integer type. Its specification defines numbers as double-precision floats (International, 2015). To isolate the impact of the representation of floats, four versions of the Hopc JavaScript compiler are tested, where each version differs only by its representation of floats, which is either tagged pointers, NaN-boxing, NuN-boxing, or self-tagging.
Hopc is an ahead-of-time JavaScript compiler that uses Scheme as its compilation backend and relies on the underlying Scheme compiler for optimizations (Serrano, 2021). This permits reusing Bigloo’s runtime to test float self-tagging in the context of JavaScript.
Figure 12 presents the results of porting Bigloo’s NaN-boxing, NuN-boxing, and self-tagging implementations to JavaScript. Performance is measured using the Hop benchmark suite, which gathers benchmarks from the popular JetStream, Octane and Sunspider JavaScript benchmark suites (Serrano, 2024). Once again, it is useful to categorize benchmarks between float-intensive benchmarks (almabench, n-body, navier-stokes, and raytrace), and benchmarks that use few or no floats.
On the float-intensive benchmarks, the relative execution times of Hopc NaN-boxing (0.26$\times$), NuN-boxing (0.26$\times$), and self-tagging (0.28$\times$) implementations are comparable, roughly 4 times faster than tagged pointers. Execution time with NaN-boxing and NuN-boxing is faster on almabench and n-body, but slower on raytrace and navier-stokes. However, on non-float benchmarks the self-tagging implementation (1.03$\times$) generally fares better due to its lower overhead than NaN-boxing (1.12$\times$) and NuN-boxing (1.16$\times$) when handling non-float objects, by up to a factor of 2.07$\times$ for crypto-sha1.
Figure 12. Execution time comparison between Hopc with 1-tag self-tagging, Hopc with NuN-boxing, and Hopc with NaN-boxing. The y-axis shows execution time relative to the original Hopc version, which uses tagged pointers. 1.0 indicates no change, lower means faster execution than the original version.
4 most 4 most
signif. signif.
expo. expo. Value
bits bits + 1 range
@@ 6.3. Application to JavaScript
7. Self-Tagging on 32-bit Architectures
Although previous sections only considered double-precision floats on 64-bit architectures, self-tagging of single-precision floats is a natural adaptation for 32-bit architectures. Indeed, the core idea, which is to superimpose a tag to a sequence of bits likely to appear in practice, can be used for any machine word size. A popular object representation on 32-bit architectures is to use 2-bit tags (4 possible tags) and heap-allocated objects aligned to 32-bit words.
The coverage can be doubled by using an adaptation of the 2-tag variant described in Section 4.4: $0.0..2.5\times 10^{-29}\cup 4.7\times 10^{-10}..8.6\times 10^{9}\cup 1.6\times 10^{29}..Infinity/Nan$ (light pink rows of Figure 13 using $tag_{1}$=00 and $tag_{2}$=11). With this variant, it would be reasonable to reserve one tag for small signed integers in the range $-2^{29}..2^{29}-1$ (-536870912 .. 536870911), and the remaining tag for heap-allocated objects, including the floats that can’t be self-tagged.
@@ 8. Related Work
Handling floats is a long-standing issue in dynamic language implementation. In the last decades, little progress has been made improving the encoding of floats, with most implementations either using heap-allocated floats or suffering from the overhead of NaN-boxing on pointers. Thus, more general strategies that use data flow analysis (Henglein, 1992; Serrano and Feeley, 1996; Petersen and Glew, 2012; Lindahl and Sagonas, 2003) and partial evaluation (Würthinger et al., 2017) have been developed to find locations where the type of a value is known across its lifecycle. This allows the generation of specialized code that safely handles fully unboxed, untagged values. Such strategies are not specific to unboxing floats. Rather, they tackle the more general problem of inferring types in dynamic languages, which allows dropping type information in either tagged pointers or NaN-boxed pointers. While more limited in scope, self-tagging solves the problem of heap-allocated floats with a new encoding that involves no program analysis.
@@
Self-tagging is straightforward to add to compilers that already use object tagging. Therefore, implementations that represent floats as tagged pointers could benefit from it with minimal implementation effort. Such popular implementations include CPython (Gross, 2023) and Google’s V8 (Sheludko and Solanes, 2020). Compilers that use NaN/NuN-boxing, such as Mozilla’s SpiderMonkey (Foundation, 2024) and Apples’s JavaScriptCore (Inc., 2024) can also benefit from the lower impact that self-tagging has on the performance of non-float types.
@@
Due to the shortcomings of tagged pointers and NaN-boxing, additional strategies are used to avoid boxing. A straightforward approach is to provide homogeneous data structures such as TypedArray in JavaScript (International, 2015) and numpy’s float64 arrays in Python (Harris et al., 2020). Type homogeneity then allows unboxing values within the collection. More generally, compilers can detect data homogeneity at run time and use context-dependent storage strategies to store data in a way that prevents pointer chasing (Bolz et al., 2013). Since self-tagging unboxes most floats in practice, it effectively prevents such pointer chasing in the case of floats without static or run time analysis.
@@
Even for general-purpose languages not specifically designed for numerical computation, ensuring reasonably fast floating-point operations is a major concern. All modern production-quality implementations surveyed for this paper devote significant effort to optimizing them. Unfortunately the academic literature is scarce when it comes to methods and techniques used by these efficient implementations.
@@
In many cases, the source code itself has to be studied to understand the implementation and compare it to self-tagging. The NuN-boxing implementation used in experiments in this paper is based on that of JavaScriptCore [^2] https://github.com/WebKit/WebKit/commit/74c9b9bbc2c7527144ac366c3f62f84aaa1a8a4e. It uses the pattern 0x0000 in the highest 16-bits to encode pointers and 0xffff to encode integers. The other patterns are used for floats. The actual types of heap-values is either encoded in the three lowest bits of the pointer itself or in a header word. For instance, as strings are frequent data type, the pattern 110 is used to encode these values and checking that a value is a string requires to check the lowest three bits and the 16 highest bits. NuN-boxing has the benefits over self-tagging that it encodes all floats uniformly and does not reserve tags for them. Its drawbacks are that it negatively impacts the type checking and accesses of all objects, it can represent integers with at most 48 bits (usually just 32 bits for practical reasons), and it can’t be used on 32-bit architectures.
@@
The technique closest to self-tagging discovered among surveyed implementations are those of CRuby [^3] https://github.com/ruby/ruby/commit/b3b5e626ad69bf22be3228f847f94e1b68f40888 and OpenSmallTalk [^4] https://clementbera.wordpress.com/2018/11/09/64-bits-immediate-floats. Both implementations use a form of self-tagging with a rotation of the mantissa close to the encoding of Section 4.3. However, bit patterns captured by both implementations exclude ±$\pm 0.0$, which is treated as a special case. This imposes complex sequences of instructions for encoding and decoding that can hardly be inlined in the generated code and risks increasing the number of branch misses as demonstrated in Section 3.4.
@@
To the best of the authors’ knowledge, this is the first peer-reviewed paper exposing the principles of self-tagging and evaluating its benefits, drawbacks, and possible variants.
@@
In this paper, self-tagging was discussed primarily in the context of dynamic languages. Yet, polymorphic languages such as OCaml and Haskell also attach information to run time values and spend considerable effort to find efficient encodings for abstract data types (Baudon et al., 2023; Kennedy and Vytiniotis, 2012). Self-tagging could be applied in these languages, either to floats or to the encoding of abstract data types.

### v3

@@ 4.4. Memory Profiling and Execution Time
Figure 7. Execution time of Bigloo with each variant of self-tagging and NaN-boxing relative to NuN-boxing on four distinct microarchitectures (from top to bottom: Amd, Intel, M2, and Risc-V). The y-axis shows execution time relative to NuN-boxing (dashed horizontal red line) on a logarithmic scale. 1.00×1.$0$ indicates no change, lower means faster, and higher means slower execution than NuN-boxing.
Execution Time Relative to NuN-Boxing (Gambit)
AMD Ryzen 7955WX, 62GB, Linux 6.12.31 x86_64, Debian 10.13, gcc 14.2.0
Intel(R) Xeon(R) W-2245, 32GB, Linux 6.12.31 x86_64, Debian 10.13, gcc 14.2.0
Apple M2 Max, 64GB, macOS 15.4.1, gcc 14.3.0
riscv sifive, u74-mc, 8GB, Linux starfive 6.1.31-starfive riscv64, gcc 14.2.0
  self-tagging (1-tag)   self-tagging (2-tag)   self-tagging (3-tag)   self-tagging (4-tag)
Figure 8. Execution time of Gambit with each variant of self-tagging relative to NuN-boxing on four distinct microarchitectures (from top to bottom: Amd, Intel, M2, and Risc-V). The y-axis shows execution time relative to NuN-boxing (dashed horizontal red line) on a logarithmic scale. 1.00×1.$0$ indicates no change, lower means faster, and higher means slower execution than NuN-boxing.
Figure 6 shows the memory allocations of Bigloo’s self-tagging variants, NaN-boxing, and NuN-boxing compared to allocated floats. Since NaN/NuN-boxing are known to allocate no floats at all, these correspond to executions where heap usage is for non-float objects only. A few self-tagging variants allocate some floats on some benchmarks (see Figure 5), but these allocations are rare enough that memory profiling of all variants is nearly identical to that of NaN/NuN-boxing.
Detailed execution time comparisons with NuN-boxing are shown in Figures 7 and 8 (Bigloo and Gambit respectively). Figure 7 also shows time for NaN-boxing, which Bigloo implements. These results are summarized in Figure 9, which provides the geometric mean of execution times relative to NuN-boxing for each float encoding, compiler, and microarchitecture, for both float-intensive and non-float benchmarks.
Summary of Execution Time Relative to NuN-Boxing
Non-float benchmarks Float benchmarks
Amd Intel M2 Risc-V Amd Intel M2 Risc-V
Bigloo 1-tag   1.01×1.01$\times$1.01 × .97×.97$\times$.97 × .97×.97$\times$.97 × .97×.97$\times$.97 × 1.20×1.20$\times$1.20 × 1.20×1.20$\times$1.20 × 1.25×1.25$\times$1.25 × 1.28×1.28$\times$1.28 ×
2-tag w/ prealloc. zero   1.01×1.01$\times$1.01 × .96×.96$\times$.96 × .98×.98$\times$.98 × .97×.97$\times$.97 × 1.22×1.$2$ 1.34×1.34$\times$1.34 × 1.33×1.$3$ 1.29×1.29$\times$1.29 ×
3-tag   1.00×1.$0$ 1.00×1.$0$ 1.00×1.$0$ 1.02×1.02$\times$1.02 × 1.13×1.13$\times$1.13 × 1.09×1.09$\times$1.09 × 1.10×1.10$\times$1.10 × 1.19×1.19$\times$1.19 ×
NaN-boxing   1.01×1.01$\times$1.01 × 1.02×1.02$\times$1.02 × 1.00×1.$0$ 1.06×1.06$\times$1.06 × .87×.87$\times$.87 × .84×.84$\times$.84 × .90×.90$\times$.90 × .90×.90$\times$.90 ×
Gambit 1-tag   .95×.95$\times$.95 × .87×.87$\times$.87 × .91×.91$\times$.91 × .88×.$\times.8$ 1.06×1.06$\times$1.06 × .94×.94$\times$.94 × 1.05×1.05$\times$1.05 × 1.03×1.03$\times$1.03 ×
w/ prealloc. zero 2-tag   .96×.96$\times$.96 × .88×.$\times.8$95×.95$\times$.95 × .89×.89$\times$.89 × $1\times1$ 1.00×1.$0$ 1.16×1.16$\times$1.16 × 1.09×1.09$\times$1.09 ×
3-tag   .96×.96$\times$.96 × .90×.90$\times$.90 × .98×.98$\times$.98 × .88×.$\times.8$ 1.09×1.09$\times$1.09 × 1.06×1.06$\times$1.06 × 1.14×1.14$\times$1.14 × 1.06×1.06$\times$1.06 ×
4-tag   .97×.97$\times$.97 × .89×.89$\times$.89 × .96×.96$\times$.96 × .89×.89$\times$.89 × .97×.97$\times$.97 × .90×.90$\times$.90 × 1.04×1.04$\times$1.04 × .94×.94$\times$.94 ×
Figure 9. Geometric mean of execution times relative to NuN-boxing of float and non-float benchmarks for each combination of float encoding, compiler, and microarchitecture. 1.00×1.$0$ indicates no change, lower means faster, and higher means slower execution than NuN-boxing.
As a general observation, none of the evaluated encodings is a better alternative overall; the best encoding depends on the nature of benchmarks, microarchitecture, and implementation. For instance with Bigloo, NaN-boxing is faster across microarchitectures on float-intensive benchmarks, but slowest on non-float benchmarks. Conversely, self-tagging generally fares better than NaN/NuN-boxing on non-float benchmarks since it introduces no overhead on non-float objects.
Bigloo’s self-tagging does not consistently match the performance of NuN-boxing on float benchmarks across microarchitectures (but is faster than allocated floats, as will be discussed in Section 4.5). The 1-tag and 3-tag variants offer the highest performance overall for Bigloo, with the 3-tag variant being faster for float benchmarks and slightly slower for non-float benchmarks. Bigloo’s 2-tag with preallocated zeros has similar average performance as 1-tag, however it shows more variability across benchmarks (see fft and pnpoly in Figure 7) since it has to test for ±0.0$\pm$ 0.0± 0.0 out-of-line. Hence, Gambit’s 2-tag variant (Section 2.5) should be preferred to that of Bigloo.
Gambit’s 4-tag variant provides consistently faster average execution times for both float and non-float benchmarks on all architectures except M2 where it is nearly the same speed as NuN-boxing. Gambit’s 1-tag variant also fares well on all benchmarks on Intel and Risc-V.
Self-tagging and NaN/NuN-boxing have competitive performance and the best choice depends on the specific setting. When floating point computation performance is important but less so than operations on other objects, self-tagging is a good option to consider. Additionally, a significant portion of this study was spent adapting existing systems (Bigloo and Gambit) to new float encodings, and self-tagging was found to be qualitatively simpler to implement as it only affects the encoding of floats, while NaN/NuN-boxing has deeper object encoding implications.
4.5. Memory Management Considerations
Execution Time Relative to Allocated Floats (Bigloo)
AMD Ryzen 7955WX, 62GB, Linux 6.12.31 x86_64, Debian 10.13, gcc 14.2.0
Intel(R) Xeon(R) W-2245, 32GB, Linux 6.12.31 x86_64, Debian 10.13, gcc 14.2.0
Apple M2 Max, 64GB, macOS 15.4.1, gcc 14.3.0
riscv sifive, u74-mc, 8GB, Linux starfive 6.1.31-starfive riscv64, gcc 14.2.0
  self-tagging (1-tag)
Figure 10. Execution time of Bigloo with self-tagging (1-tag) relative to allocated floats on four distinct microarchitectures (from top to bottom: Amd, Intel, M2, and Risc-V). The y-axis shows execution time relative to the version that heap allocates all floats (dashed horizontal red line) on a logarithmic scale. 1.00×1.$0$ indicates no change, lower means faster execution than allocated floats.
Execution Time Relative to Allocated Floats (Gambit)
AMD Ryzen 7955WX, 62GB, Linux 6.12.31 x86_64, Debian 10.13, gcc 14.2.0
Intel(R) Xeon(R) W-2245, 32GB, Linux 6.12.31 x86_64, Debian 10.13, gcc 14.2.0
Apple M2 Max, 64GB, macOS 15.4.1, gcc 14.3.0
riscv sifive, u74-mc, 8GB, Linux starfive 6.1.31-starfive riscv64, gcc 14.2.0
  self-tagging (4-tag)
Figure 11. Execution time of Gambit with self-tagging (4-tag) relative to allocated floats on four distinct microarchitectures (from top to bottom: Amd, Intel, M2, and Risc-V). The y-axis shows execution time relative to the version that heap allocates all floats (dashed horizontal red line) on a logarithmic scale. 1.00×1.$0$ indicates no change, lower means faster execution than allocated floats.
Execution Time with Non-Empty Heap
AMD Ryzen 7955WX, 62GB, Linux 6.12.31 x86_64, Debian 10.13, gcc 14.2.0
Execution time (s) fft Refer to caption fibfp Refer to caption mbrot Refer to caption nucleic Refer to caption pnpoly Refer to caption ray Refer to caption simplex Refer to caption sum1 Refer to caption sumfp Refer to caption Preallocated heap data size (bytes) Refer to caption
Figure 12. Execution time of float benchmarks with extra data allocated on heap before execution with Bigloo and Gambit on Amd. The x-axis shows the size of preallocated data in bytes on a logarithmic scale. The y-axis represents execution time in seconds (mean of 10 repetitions). Execution time without preallocated data is added at $x=0$. Dashed lines correspond to times with allocated floats. Solid lines are with self-tagging.
Figures 10 and 11 show execution times of the best self-tagging variants on each implementation (Bigloo 1-tag, and Gambit 4-tag respectively) relative to allocated floats. For Bigloo, decreased memory allocations correlates with lower execution time on all architectures (see Figure 6) with the exception of sum1, an I/O bound benchmark.
Conversely, Gambit execution time slightly increases on many float benchmarks. This difference is caused by its use of a stop-and-copy garbage collector with bump allocation, while Bigloo uses the Boehm mark-and-sweep GC (Boehm and Weiser, 1988). For programs with a nearly empty heap (default of R7RS benchmarks), the overhead of self-tagging’s encoding/decoding outweighs that of bump allocation.
However, since self-tagging improves performance by allocating fewer floats, its impact is more pronounced for programs that spend more time in memory management tasks. This section presents an experiment to measure execution time in a more realistic setting where benchmarks are executed with live data on the heap, thus putting more pressure on the garbage collector.
For this experiment, the R7RS benchmark suite is modified to allocate a vector before the execution of each benchmark. The size of the allocated vector is given as a parameter of the benchmark. Each float benchmark is then executed with vectors of $10^{4}$ to $10^{8}$ fields, which correspond to about 80 kB and 800 MB respectively. An execution is also done with no preallocated vector.
Figure 12 compares execution times for Bigloo and Gambit with self-tagging (1-tag and 4-tag respectively) and allocated floats with increasing vector sizes. With a nearly empty heap, and thus little strain on the garbage collector, the bump allocator of Gambit with allocated floats sometimes outperforms self-tagging since it does not incur the cost of encoding and decoding self-tagged floats. Once the heap is preloaded with about 1 MB (which is the size of the L2 cache per core on the Amd machine) the strain of the garbage collector starts to outweigh the cost of encoding and decoding self-tagged floats, thus self-tagging hereafter outperforms allocated floats.
4.6. Branch Prediction Considerations
This section discusses the impact of branch misprediction on the performance of self-tagging by presenting another variant of self-tagging that was developed based on intuitions that turned out to be wrong. This negative result is detailed here to guide implementers when adapting self-tagging.
Self-tagging variants from previous sections use high exponent bits of floats. This is convenient since it captures contiguous ranges of floats at the cost of a small encoding/decoding overhead. Alternatively, low bits of a float’s mantissa can be used. For instance, by reserving the tag 000 for self-tagged floats and heap allocating all floats whose low mantissa bits are not 000.
Execution Time of Self-Tagging with Low Bits (Bigloo)
AMD Ryzen 7955WX, 62GB, Linux 6.12.31 x86_64, Debian 10.13, gcc 14.2.0
  self-tagging (2-tag, mantissa low bits)
Figure 13. Execution time of Bigloo with self-tagging using the mantissa low bits relative to allocated floats on the Amd microarchitecture. The y-axis shows execution time relative to the version that heap allocates all floats (dashed horizontal red line) on a logarithmic scale. 1.00×1.$0$ indicates no change, lower means faster, and higher means slower execution than allocated floats.
This encoding has no overhead since a self-tagged float corresponds exactly to its IEEE754 value. It comes as the cost of no longer capturing contiguous ranges of floats, but rather a set uniformly spread across all possible floats. As in previous techniques, the proportion of captured floats can be increased by assigning more tags to self-tagging. This variant is implemented in Bigloo with tags 000 and 100, thus self-tagging 1/41/41 / 4 of all floats, in particular all floats $n$ that are integers $|n|\leq 2^{51}$.
Figure 13 compares execution times of this variant to allocated floats. As expected, benchmarks that extensively use floats corresponding to integers (fibfp and sumfp) execute faster, even faster than previous variants (see Figure 10), due to most floats avoiding heap allocation. However, other benchmarks (mbrot and nucleic) are slower despite allocating about 1/41/41 / 4 fewer floats.
This unpredictability is explainable by profiling the number of branch mispredictions caused by this variant. Figure 14 shows missed branch predictions of float encodings, including self-tagging with mantissa bits. This encoding causes a steep increase in missed branch predictions on some benchmarks, namely mbrot and nucleic. This stems from the fact that when tags occupy the low bits of the mantissa, even the smallest variations cause a float’s representation to switch between self-tagged and tagged pointer in a hard to predict way. This is beyond the capabilities of the branch predictor and execution suffers costly misprediction penalties (Bryant and O’Hallaron, 2016).
Branch Misprediction of Self-Tagging with Low Bits (Bigloo)
AMD Ryzen 7955WX, 62GB, Linux 6.12.31 x86_64, Debian 10.13, gcc 14.2.0
  self-tagging (2-tag, mantissa low-bits)   self-tagging (1-tag)
  self-tagging (2-tag w/ prealloc. zero)   NaN-boxing   NuN-boxing
Figure 14. Branch mispredictions of Bigloo with each variant of self-tagging, NaN-boxing, and NuN-boxing relative to allocated floats by the Amd microarchitecture. The y-axis shows branch mispredictions relative to allocated floats (dashed horizontal red line) on a logarithmic scale. 1.00×1.$0$ indicates no change, lower means less, and higher means more mispredictions.
This result refutes the intuition that self-tagging with mantissa bits may be a good compromise due to its very cheap encoding/decoding. It also hints at the necessity to avoid introducing tests for special values in encoding/decoding, such as the test for ±0.0$\pm$ 0.0± 0.0 in the 2-tag self-tagging variant with preallocated zeros, which has a higher occurrence of branch mispredictions than 1-tag self-tagging in Figure 14.
5. Self-Tagging on 32-bit Architectures
Although previous sections only considered double-precision floats on 64-bit architectures, self-tagging of single-precision floats is a natural adaptation for 32-bit architectures. Indeed, the core idea, which is to superimpose a tag to a sequence of bits likely to appear in practice, can be used for any machine word size. A popular object representation on 32-bit architectures is to use 2-bit tags (4 possible tags) and heap allocated objects aligned to 32-bit words.
The coverage can be doubled by using an adaptation of the 2-tag variant described in Section 2.5: $0.0..2.5\times 10^{-29}\cup 4.7\times 10^{-10}..8.6\times 10^{9}\cup 1.6\times 10^{29}..Infinity/NaN$ (light pink rows of Figure 15 using $tag_{1}$=00 and $tag_{2}$=11). With this variant, it would be reasonable to reserve one tag for small signed integers in the range $-2^{29}..2^{29}-1$ (-536870912 .. 536870911), and the remaining tag for heap allocated objects, including the floats that can’t be self-tagged.
4 most signif. 4 most signif.
expo. bits expo. bits + 1 Value
@@ 5. Self-Tagging on 32-bit Architectures
4 most signif. 4 most signif.
expo. bits expo. bits + 1 Value
@@ 6. Related Work
Handling floats is a long-standing issue in dynamic language implementation. In the last decades, little progress has been made improving the encoding of floats, with most implementations either using heap allocated floats or suffering from the overhead of NaN-boxing on pointers. Thus, more general strategies that use data flow analysis (Henglein, 1992; Serrano and Feeley, 1996; Petersen and Glew, 2012; Lindahl and Sagonas, 2003; Kawai, 2008) and partial evaluation (Würthinger et al., 2017) have been developed to find locations where the type of a value is known across its lifecycle. This allows the generation of specialized code that safely handles fully unboxed, untagged values. Such strategies are not specific to unboxing floats. Rather, they tackle the more general problem of inferring types in dynamic languages, which allows dropping type information in either tagged pointers or NaN-boxed pointers. While more limited in scope, self-tagging solves the problem of heap allocated floats with a new encoding that involves no program analysis.
@@
Self-tagging is straightforward to add to compilers that already use object tagging. Therefore, implementations that represent floats as tagged pointers could benefit from it with minimal implementation effort. Such popular implementations include CPython (Gross, 2023) and Google’s V8 (Sheludko and Solanes, 2020). Compilers that use NaN/NuN-boxing, such as Mozilla’s SpiderMonkey (Foundation, 2024) and Apples’s JSC (Inc., 2024) can also benefit from the lower impact that self-tagging has on the performance of non-float types.
@@
Due to the shortcomings of tagged pointers and NaN-boxing, additional strategies are used to avoid boxing. A straightforward approach is to provide homogeneous data structures such as TypedArray in JavaScript (International, 2015) and numpy’s float64 arrays in Python (Harris et al., 2020). Type homogeneity then allows unboxing values within the collection. More generally, compilers can detect data homogeneity at run time and use context-dependent storage strategies to store data in a way that prevents pointer chasing (Bolz et al., 2013). Since self-tagging avoids heap allocating most floats in practice, it effectively prevents such pointer chasing in the case of floats without static or run time analysis.
@@
Even for general-purpose languages not specifically designed for numerical computation, ensuring reasonably fast float operations is a major concern. All production-quality implementations surveyed for this paper devote significant efforts to optimizing them. Unfortunately, the academic literature is scarce when it comes to techniques used by these efficient implementations.
@@
In many cases, the source code itself has to be studied to understand the implementation and compare it to self-tagging. The NuN-boxing implementation used in experiments in this paper is based on that of JSC [^2] https://github.com/WebKit/WebKit/commit/74c9b9bbc2c7527144ac366c3f62f84aaa1a8a4e. NuN-boxing has the benefits over self-tagging that it encodes all floats uniformly and does not reserve tags for them. Its drawbacks are that it negatively impacts the type checking and accesses of all objects, it can represent integers with at most 48 bits (usually 32 bits for practical reasons), and it can’t be used on 32-bit architectures.
NaN/NuN-boxing are trade-offs between the performance of floats and pointers (NaN-boxing favors floats, NuN-boxing favors pointers). Self-tagging offers a different trade-off between the performance of frequently used floats and other floats. The fact that some ranges of floats are more commonly used has been studied in scientific computing where real world data typically fall into a narrow range of values (Brown et al., 2007; Lindstrom and Isenburg, 2006; Afroozeh et al., 2023; Zurstrassen et al., 2023). This is in accordance with the results presented in Figure 5.
@@
The techniques closest to self-tagging discovered among surveyed implementations are those of CRuby [^3] https://github.com/ruby/ruby/commit/b3b5e626ad69bf22be3228f847f94e1b68f40888, OpenSmallTalk [^4] https://clementbera.wordpress.com/2018/11/09/64-bits-immediate-floats and the MonNom calculus (Muehlboeck and Tate, 2021). CRuby and OpenSmallTalk use a variant close to the 1-tag encoding of Section 2.4. The MonNom implementation from (Muehlboeck and Tate, 2021) uses a variant of the 2-tag encoding of Section 2.3 with 2-bit tags instead of 3-bit. Bit patterns captured by these implementations all exclude ±0.0$\pm$ 0.0± 0.0, which is treated as a special case. This imposes complex sequences of instructions for encoding/decoding that can hardly be inlined and risks increasing the number of branch misses as discussed in Section 4.6.
@@
To the best of the authors’ knowledge, this is the first work exposing the principles of self-tagging, carefully implementing it in optimizing compilers, and evaluating its benefits, drawbacks, and multiple variants.
@@
In this paper, self-tagging was discussed in the context of dynamic languages. Yet, polymorphic languages such as OCaml and Haskell also attach information to run time values and spend considerable effort to find efficient encodings for abstract data types (Baudon et al., 2023; Kennedy and Vytiniotis, 2012). Self-tagging could be applied in these languages, either to floats or to the encoding of abstract data types.